OpenStudy (anonymous):

The price of 9-volt batteries is increasing according to the function below, where t is years after January 1, 1980. During what year will the price reach $5? P(t) = 1.1 e0.047t

OpenStudy (anonymous):

P(t) = 1.1 e^0.047t

OpenStudy (anonymous):

Substitute P(t)=5 and solve for t

OpenStudy (anonymous):

so the setup of it would be 5=1.1*e^o.o47t? but im not good at exponents and have no idea how to figure this out

OpenStudy (anonymous):

Yeah. So first u bring the 1.1 over to the left hand side . 5/1.1= e^0.047t Apply natural logarithm to both side you get ln (5/1.1) = ln e^0.047t Bring the power of ln e^0.047t to the front ln (5/1.1)= 0.047t ln e Ln e = 1 ln (5/1.1)=0.047t t= (ln(5/1.1))/0.047 =32.2

OpenStudy (anonymous):

oh i see now.. thanks. imma write this down on some paper so i have some reference in case i come to another problem like this. thanks again

OpenStudy (anonymous):

no problem :D